A company reports the following​ balances:

​12/31/2017

​12/31/2018

Common Stock

​$14,000

​$165,000

Paidminus−in

Capital in Excess of Par

​10,000

​10,000

Land

​93,000

​109,000

Retained Earnings

​60,000

​76,000

Treasury Stock

​16,000

​18,000

During​ 2018, net income of​ $29,000 was reported. No treasury stock was sold during 2018. No common stock was retired during 2018. No land was sold during 2018. What financing activities are reported on the statement of cash flows prepared with the indirect method for the year ending December​ 31, 2018?

A.

sale of treasury stock​ $2,000 and Payment of dividends​ $13,000

B.

payment of dividends​ $16,000, Sale of common stock​ $167,000, and Purchase of treasury stock​ $2,000

C.

sale of treasury stock​ $2,000, purchase of land​ $10,000.

D.

purchase of treasury stock​ $2,000, Payment of dividends​ $13,000 and Sale of common stock​ $151,000
Answer #1

Answer:

Correct answer is:

D. purchase of treasury stock​ $2,000, Payment of dividends​ $13,000 and Sale of common stock​ $151,000

Explanation:

Sale of common stock = ​$165,000 - $14000 = $151,000

Dividend payment = Retained earning beginning balance + Net Profit - Retained earning ending balance

= 60000 + 29000 - 76000

= $13,000

Purchase of treasury stock = Ending balance treasury stock - Beginning balance treasury stock = 18000 - 16000 = $2,000

Purchase of land is not reported under financing activities

As such option D is correct and other options A, B and C are incorrect.